Downtime due to power outage


Milou, Tintin, and Fysast-1 are back in production. Bianca is back in test production. Still working on Smog.

There was a power outage this morning. Our systems at UPPMAX were affected by this. We are investigating the extent, to see what systems are affected, and if any permanent damage was done. All our efforts are focused on getting all clusters back up again.

Update 10:00

There is a problem accessing several volumes on Pica. We are in contact with the manufacturer. Until Pica is available, compute nodes are drained on Tintin, Milou and Fysast1, meaning no new jobs can start. Login nodes are up, but home directories are unavailable.

Most of Irma is in production. Some compute nodes need further investigation.

Bianca is not yet in production. We need to power up storage system Castor first (it was shut down to stop the rapid warming of our computer room). This will probably be done in the afternoon.

Update 11:00

Pica is alive and well. Milou, Tintin, and Fysast1 are back in production.

We are still working on Bianca.

Update 14:10

Still working on Bianca and Smog. We are expecting Bianca to be up later today.

Update 16:00

Biance is back in test production, which means test projects on Bianca can run just as before.

Old System News